OpenStudy (firecub17):

Use the function below to find F(3) F(x)=(1/4)^x

OpenStudy (3mar):

If I may help?

OpenStudy (firecub17):

Please do! @3mar

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (3mar):

Thanks. The idea is to plug in the value of x in the function, i.e replace x with its value \[\large f(x)=[\frac{ 1 }{ 4 }]^x\] \[f(3)=[\frac{ 1 }{ 4 }]^3=[\frac{ 1 }{ 4^3 }]=\frac{ 1 }{ 64 }\]

OpenStudy (firecub17):

Thank you!! @3mar

OpenStudy (3mar):

You are welcome! Any Help... Any Time... + Thank you for the medal!